OpenStudy (anonymous):

How do you find the hypotenuse of a right triangle when you are given the lengths of the legs? A=9 B=12 C=?

OpenStudy (anonymous):

Here is the equation a^2 + b^2 = c^2

OpenStudy (anonymous):

(9x9) + (12x12) = (c x c) 81 + 144 = c ^2 What is the square root of 225?

OpenStudy (anonymous):

15. Thank you!